OpenStudy (anonymous):

Can someone help me with this? You go to school in a college town. You know that there are 2000 students enrolled in the school, but you don’t know the population of the town (without students). You walk up and down the main streets of the town, stop people, and ask them if they are students or not. You ask 100 people, and 60 of them say they are students. Estimate the nonstudent population of the town.

OpenStudy (anonymous):

this one?

Directrix (directrix):

Let's see what we can do. @cnr

OpenStudy (anonymous):

I already figured out one of the others I asked about

Directrix (directrix):

Great.

OpenStudy (anonymous):

my computer is freezing so I will have to get back on it a little while!! So frustrating. Thank you for your help. I sincerely appreciate it.

Directrix (directrix):

This is a sampling problem. Let me think aloud. 100 people sampled, 60 are students so 40 are non-students Ratio of students to non-students in the town is 60/40 Assume people on the street are randomly mixed students and non-students.

Directrix (directrix):

Does that make sense? Now, we have to craft a proportion of some sort.

OpenStudy (anonymous):

yes it does, that is about as far as I got

Directrix (directrix):

60/40 = 2000/NS where NS is non-students. If the ratio of students to NS in the sample is 60/40, then 60 to 40 should be the ratio of the 2000 students in town to all of the NS in the town.

OpenStudy (anonymous):

my guess was 1500

Directrix (directrix):

Solve 60/40 = 2000/NS and see what NS is. Let's see if it makes sense as an answer.

OpenStudy (anonymous):

I wasn't understanding my math when I used that answer.

OpenStudy (anonymous):

My computer is getting re booted, hopefully you will be on when I get back ;) thanks again if not

Directrix (directrix):

This site goes up and down like a yo-yo sometimes. I'm here and thinking that this equation does not work. I need to think while you re-boot. @cnr

Directrix (directrix):

@cnr

Directrix (directrix):

60/40 = 2000/NS Let x = NS (Non-students in town) 3/2 = 2000/x x = 1333 approx --> non-student town population

Directrix (directrix):

For every 3 students, there are 2 non-students. @cnr